please help me with this problem.
if f(x)= 2x
find f(x+3)

Answer by nerdybill(7384)

if f(x)= 2x
find f(x+3)
.
simply replace x in:
f(x)= 2x
with "x+3"
.
f(x+3)= 2(x+3)
f(x+3)= 2x+6 (this is what they're looking for)